# Graph-Reward-SQL: Execution-Free Reinforcement Learning for Text-to-SQL via Graph Matching and Stepwise Reward

### v2

@@ Abstract
Reinforcement learning (RL) has been widely adopted to enhance the performance of large language models (LLMs) on Text-to-SQL tasks. However, existing methods often rely on execution-based or LLM-based Bradley–Terry reward models (BTRMs). The former suffers from high execution latency caused by repeated database calls, whereas the latter imposes substantial GPU memory overhead, both of which significantly hinder the efficiency and scalability of RL pipelines. To this end, we propose a novel Text-to-SQL RL fine-tuning framework named Graph-Reward-SQL, which employs the GMNScore outcome reward model. We leverage SQL graph representations to provide accurate reward signals while significantly reducing inference time and GPU memory usage. Building on this foundation, we further introduce StepRTM, a stepwise reward model that provides intermediate supervision over Common Table Expression (CTE) subqueries. This encourages both functional correctness and structural clarity of SQL. Extensive comparative and ablation experiments on standard benchmarks, including Spider and BIRD, demonstrate that our method consistently outperforms existing reward models.
@@
Han Weng, Puzhen Wu, Cui Longjie, Yi Zhan, Boyi Liu, Yuanfeng Song, Dun Zeng, Yingxiang Yang, Qianru Zhang, Dong Huang, Xiaoming Yin, Yang Sun, Xing Chen ByteDance Inc.
@@ 1 Introduction
Text-to-SQL (Tai et al., 2023; Li et al., 2024b; Shi et al., 2025) aims to translate natural language into structured database queries, playing a crucial role in democratizing data access by enabling non-technical users to interact effectively with relational databases. A significant amount of work is devoted to the fine-tuning of a foundational model, where Reinforcement Learning (RL) has recently been shown to effectively enhance model performance (Pourreza et al., 2025b; Berdnyk and Collery, 2025; Ma et al., 2025). Among these efforts, the careful design of the Reward Model (RM) is a crucial challenge, as the quality of the reward signal directly influences policy optimization during fine-tuning.
@@
In RL-based Text-to-SQL approaches, execution accuracy remains a dominant signal Nguyen et al. (2025); Ma et al. (2025); Pourreza et al. (2025b); Berdnyk and Collery (2025), providing intuitive feedback based on query correctness. Additionally, the LLM-based Bradley–Terry reward model (BTRM) Christiano et al. (2017) has been adapted for code generation by deriving preference pairs from execution outcomes Zeng et al. (2025a). Structural rewards based on abstract syntax tree (AST) have also been explored to capture syntactic similarity Shojaee et al. (2023). However, each approach has significant limitations in the Text-to-SQL task. Execution-based rewards introduce significant latency due to runtime database access. LLM-based BTRM incurs high computational and memory costs, limiting scalability. AST matching-based similarity is prone to false negatives, where syntactically divergent queries that are semantically equivalent are penalized, leading to inaccurate reward signals. These limitations underscore a key challenge in Text-to-SQL RL: designing an efficient reward model that can replace execution-based signals without compromising performance.
@@
To address the above limitations, we introduce Graph-Reward-SQL, a novel RL framework for Text-to-SQL tasks. This framework incorporates two complementary reward models: Graph Matching Network Score (GMNScore) and Stepwise Relational Operator Tree Match (StepRTM). GMNScore serves as an outcome-based reward, which evaluates the generated SQL queries using the Graph Matching Network (GMN) without requiring execution. GMN utilizes learned graph embeddings to assess functional equivalence, capturing the deep semantics of SQL queries Zhan et al. (2025). In contrast to execution-based rewards, GMNScore eliminates the need for costly database executions, resulting in a significant speed-up. Furthermore, compared to LLM-based Bradley-Terry reward models (BTRM), GMNScore substantially reduces GPU memory consumption due to the lightweight architecture of GMN. Additionally, StepRTM provides intermediate feedback through a stepwise reward mechanism that evaluates the generation of Common Table Expression (CTE) subqueries, complementing GMNScore.
@@
The above design offers three notable advantages. (i) Superior Training Efficiency: Our method significantly reduces time cost and GPU memory usage compared to existing outcome reward models, leading to enhanced overall training efficiency for reinforcement learning. (ii) Intermediate Feedback Integration: Unlike existing models that focus solely on outcome evaluation, our reward model incorporates intermediate feedback by leveraging the structure of CTE SQL. This provides richer feedback during training, improving performance. (iii) Strong Empirical Performance: Extensive ablation studies and evaluations on the Spider Yu et al. (2018) and BIRD Li et al. (2024b) Text-to-SQL benchmarks validate the superiority of our reward model. The results consistently demonstrate that our approach outperforms multiple strong reward model baselines, highlighting its effectiveness.
Our main contributions can be summarized as follows:
@@
  • We propose GMNScore, the first outcome reward model that leverages GMN to replace execution-based rewards, achieving both higher efficiency and better performance.
@@
  • We design a novel stepwise reward model StepRTM, which utilizes CTE SQL to deliver stepwise supervision by matching each subquery, resulting in improved performance.
@@
  • Extensive experiments show that our reward models consistently improve performance while maintaining high inference efficiency and low memory consumption.
@@ 2 Related Work
Text-to-SQL. Text-to-SQL is a key task in Natural Language Processing (NLP) that involves transforming queries expressed in natural language into executable SQL queries (Tai et al., 2023; Li et al., 2024b; Shi et al., 2025). With the increasing deployment of large language models (LLMs), agentic frameworks (Wang et al., 2025; Pourreza et al., 2025a; Lei et al., 2024) have been introduced to enhance Text-to-SQL tasks. These frameworks enable LLMs to interact with databases through iterative reasoning and external tools. Code Foundation Models such as DeepSeek-Coder (Guo et al., 2024) and Qwen2.5-Coder (Hui et al., 2024) provide the backbone for these agentic systems, enabling structured reasoning and code generation. Several approaches aim to improve LLM performance in Text-to-SQL tasks, including direct fine-tuning (Li et al., 2024a; Yang et al., 2024; Pourreza and Rafiei, 2024), as well as techniques such as prompt design (Pourreza and Rafiei, 2023; Dong et al., 2023; Gao et al., 2024) and schema linking (Guo et al., 2019; Wang et al., 2020; Lei et al., 2020; Lee et al., 2025) to further optimize results.
@@
Reinforcement Learning and Reward Model. RL has become a important paradigm for effectively fine-tuning Code Foundation Models. Policy optimization methods, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024), have been explored. However, the effectiveness of RL training heavily relies on the quality of reward signals, making the design of reward models a critical aspect (Trella et al., 2023). Several contributions to RL-based code generation have advanced reward model strategies. Notable works include CodeRL (Le et al., 2022), which leverages execution feedback, PPOCoder (Shojaee et al., 2023), which integrates semantic matching of abstract syntax trees, and AceCoder (Zeng et al., 2025a), which applies an LLM-based Bradley-Terry Reward Model.
@@
The execution-based reward model for Text-to-SQL was initially introduced by Zhong et al. (2017). Recent advancements have introduced continuous reward scores based on keyword matching (Nguyen et al., 2025) and leveraged LLMs to generate and iteratively refine reward model design (Berdnyk and Collery, 2025). Alongside these developments, reasoning models such as DeepSeek-R1 (Guo et al., 2025) have advanced RL in reasoning tasks, leading to the introduction of more sophisticated reward model designs. For example, SQL-R1 (Ma et al., 2025) incorporates format and length constraints, while Reasoning-SQL (Pourreza et al., 2025b) employs more complex reward structures, such as schema linking feedback, n-gram similarity scores, and LLM-based judgment. Despite these enhancements, execution-based rewards continue to play a central role in above RL-based Text-to-SQL approaches.
@@
Current methods overlook the computational overhead of execution-based and LLM-based reward models and fail to fully exploit the deep semantic structure of SQL queries. Additionally, these approaches focus solely on evaluating the final generated SQL, neglecting the potential of leveraging intermediate supervision signals throughout the SQL generation process. To address these issues, we propose an execution-free outcome reward model and a stepwise reward mechanism. These methods significantly reduce computational overhead while providing more effective reward signals for RL-based Text-to-SQL task.
@@ 3.1 Problem Formulation
In the standard Text-to-SQL setting, $x$ denotes a natural language query, $\hat{q}$ and $q^{\star}$ denote the generated SQL query and reference SQL, respectively. In this work, we mainly use Proximal Policy Optimization (PPO) (Schulman et al., 2017), which optimizes the policy model $\pi_{\theta}$ by maximizing:
@@ 3.2 Summary of Existing Reward Models
Recognizing the great importance of reward models in RL, we discuss three types of main reward models. As summarized in Table 1, we compare these models with our proposed reward models in terms of time cost and GPU memory usage during inference. Additionally, the final performance of all reward models is evaluated and ranked, as described in Section 6.1. Detailed information on these comparisons can be found in the Appendix F.
@@
EX (Pourreza et al., 2025b) Execution Outcome Slow N/A 3
BTRM (Zeng et al., 2025a) LLM Outcome Moderate High 5
@@
Table 1: Comparison of Reward Models in RL for Text-to-SQL Tasks. Our proposed GMNScore and StepRTM achieves better performance while significantly reducing time and memory costs.
@@
For the Text-to-SQL task, the execution accuracy serves as the most direct reward signal, providing a discrete score based on whether the generated SQL query yields the correct result upon execution. We use a discrete reward model with finer-grained feedback based on syntax error Pourreza et al. (2025b) and runtime diagnostics following Shojaee et al. (2023). Given a generated SQL $\hat{q}$ and reference SQL $q^{\star}$, the formulation is listed as:
@@
However, the EX has notable limitations. When the database contains poor quality data (e.g., limited, missing, or inconsistent entries) or structural issues (e.g., redundancy or anomalies), different queries may produce identical results (Zhong et al., 2020). The Test Suite (TS) (Zhong et al., 2020) attempted to address this issue, but as shown in Zhan et al. (2025), false positives and false negatives remain unavoidable. Additionally, repeatedly executing SQL queries introduces significant computational overhead, increasing training time. More details about EX are provided in the Appendix E.
@@
Matching based Reward.
In (Nguyen et al., 2025), keyword matching is used for SQL queries, while n-gram similarity is used in Reasoning-SQL (Pourreza et al., 2025b) to capture overlapping token sequences. Matching-based methods are fast but may assign negative rewards to semantically equivalent SQL queries that differ in syntax, which should ideally be considered correct. In broader code generation tasks, the PPOCoder (Shojaee et al., 2023) uses semantic matching of abstract syntax trees and data flow graphs. However, it still focuses on surface-level structure and does not fully capture the deep semantic information.
@@ 4 Methodology
We introduce Graph-Reward-SQL, a novel framework designed to enhance SQL generation through two key innovations. First, we propose GMNScore, which replaces EX, reducing time costs while maintaining the accuracy of reward signals without requiring database execution. Second, we introduce StepRTM, a stepwise reward model based on the Relational Operator Tree (ROT) representation of CTE SQL, which provides intermediate feedback.
Figure 1: Graph-Reward-SQL employs PPO, where rewards drive policy updates. An example illustrates a limitation of EX: the generated query ’WHERE age > 34’ and the reference ’WHERE age >= 34’ receive identical scores despite their semantic difference. In contrast, our framework uses GMNScore to evaluate deep semantic similarity. Additionally, we introduce StepRTM, a stepwise reward model tailored for CTE-SQL to address the lack of intermediate rewards. A mock stepwise score is provided for illustration; see Figure 2 for more details.
@@ 4.1 Relational Operator Tree (ROT)
To address this gap, we leverage the Relational Operator Tree (ROT) to represents SQL queries as trees of relational algebra operators. Each node in the tree corresponds to a specific logical operation (e.g., Join, Project, Filter), while the tree structure itself reflects the dependencies and execution order of the query. In practice, we use Apache Calcite (Begoli et al., 2018) to generate ROTs, which compiles SQL into a canonical intermediate representation called RelNode. This format includes various optimizations, such as operator reordering and clause simplification, resulting in normalized logical plans that are more resilient to surface-level differences. Similarly to CFGs and DFGs, the RelNode format can also integrate control dependencies and data flow as edges (Zhan et al., 2025). This enables the creation of more comprehensive graph representations that are essential for deeper query analysis.
@@ 4.2 FuncEvalGMN
After obtaining the SQL graph representations, we employ a Graph Matching Network (GMN) (Li et al., 2019) trained on SQL pairs (Zhan et al., 2025) to assess functional equivalence. The model introduces global positional encoding and a cross-graph attention mechanism. It is trained using contrastive learning for pretraining and supervised learning to capture deep semantic similarity of SQL queries. The similarity between two queries is computed as the negative Euclidean distance between their final graph-level embeddings: $s(h_{G_{1}},h_{G_{2}})=\big{\lVert}h_{G_{1}}-h_{G_{2}}\big{\rVert}_{2},$ where $h_{G_{1}}$ and $h_{G_{2}}$ are computed by the GMN, considering the joint representations of $G_{1}$ and $G_{2}$. This approach, first introduced in FuncEvalGMN (Zhan et al., 2025), is described in further detail in Appendix M.

### v3

@@ Abstract
Reinforcement learning (RL) has been widely adopted to enhance the performance of large language models (LLMs) on Text-to-SQL tasks. However, existing methods often rely on execution-based or LLM-based Bradley–Terry reward models. The former suffers from high execution latency caused by repeated database calls, whereas the latter imposes substantial GPU memory overhead, both of which significantly hinder the efficiency and scalability of RL pipelines. To this end, we propose a novel reward model framework for RL-based Text-to-SQL named Graph-Reward-SQL, which employs the GMNScore outcome reward model. We leverage SQL graph representations to provide accurate reward signals while significantly reducing time cost and GPU memory usage. Building on this foundation, we further introduce StepRTM, a stepwise reward model that provides intermediate supervision over Common Table Expression (CTE) subqueries. This encourages both functional correctness and readability of SQL. Extensive comparative and ablation experiments on standard benchmarks, including Spider and BIRD, demonstrate that our method consistently outperforms existing reward models.
@@
Han Weng1,2thanks: Work was done during the internship at ByteDance., Puzhen Wu2, Longjie Cui211footnotemark: 1, Yi Zhan211footnotemark: 1, Boyi Liu2, Yuanfeng Song2, Dun Zeng2, Yingxiang Yang2, Qianru Zhang2, Dong Huang2,322footnotemark: 2, Xiaoming Yin2, Yang Sun2thanks: Corresponding authors , Xing Chen2 1 Beijing University of Posts and Telecommunications, China 2 ByteDance, China 3 Institute of Data Science, National University of Singapore
@@ 1 Introduction
Text-to-SQL (Tai et al., 2023; Li et al., 2024b; Shi et al., 2025) aims to translate natural language into structured database queries and plays a crucial role in democratizing data access by enabling non-technical users to interact with relational databases more effectively. A significant body of work has focused on fine-tuning foundational models, with recent studies showing that Reinforcement Learning (RL) can effectively enhance model performance (Pourreza et al., 2025b; Berdnyk and Collery, 2025; Ma et al., 2025). Among these efforts, the careful design of the reward model is a crucial challenge, as the quality of the reward signal directly influences policy optimization during fine-tuning.
@@
In RL-based Text-to-SQL approaches, execution accuracy remains a dominant signal Nguyen et al. (2025); Ma et al. (2025); Pourreza et al. (2025b); Berdnyk and Collery (2025), providing intuitive feedback based on query correctness. Additionally, the Bradley–Terry reward model (BTRM) Christiano et al. (2017) has been adapted for code generation by deriving preference pairs from execution outcomes Zeng et al. (2025a). Structural rewards based on abstract syntax tree (AST) have also been explored to capture syntactic similarity Shojaee et al. (2023). However, each approach has significant limitations in the Text-to-SQL tasks. Execution-based rewards introduce significant latency due to runtime database access. The LLM-based BTRM incurs high computational and GPU memory costs, which limits its scalability. AST matching-based similarity is prone to false negatives, where syntactically divergent queries that are semantically equivalent are penalized, leading to inaccurate reward signals. These limitations point to a key challenge in RL-based Text-to-SQL: designing an efficient reward model that can replace execution-based signals without compromising performance.
@@
To address the above limitations, we introduce Graph-Reward-SQL, a novel reward model framework for RL-based Text-to-SQL. This framework incorporates two complementary reward models: Graph Matching Network Score (GMNScore) and Stepwise Relational Operator Tree Match (StepRTM). GMNScore serves as an outcome reward model, which evaluates the generated SQL queries using the Graph Matching Network (GMN) without requiring execution. GMN utilizes learned graph embeddings to assess functional equivalence, capturing the deep semantics of SQL queries Zhan et al. (2025). In contrast to execution-based rewards, GMNScore eliminates the need for costly database executions, resulting in a significant speed-up. Furthermore, compared to LLM-based Bradley-Terry reward models (BTRM), GMNScore substantially reduces GPU memory consumption due to the lightweight architecture of GMN. Additionally, StepRTM provides intermediate feedback through a stepwise reward mechanism that evaluates the generation of Common Table Expression (CTE) subqueries, complementing GMNScore.
@@
The above design offers three notable advantages. (i) Superior Training Efficiency: Our method significantly reduces time cost and GPU memory usage compared to existing outcome reward models, leading to enhanced overall training efficiency for RL. (ii) Intermediate Feedback Integration: Unlike existing reward models that focus solely on outcome result, our framework incorporates intermediate evaluation by leveraging the structure of CTE SQL. This provides richer feedback during training, improving performance and readability. (iii) Strong Empirical Performance: Extensive ablation studies and evaluations on the Spider Yu et al. (2018) and BIRD Li et al. (2024b) Text-to-SQL benchmarks validate the superiority of our reward models. The results consistently demonstrate that our approach outperforms multiple strong reward model baselines, highlighting its effectiveness.
Our contributions are summarized as follows:
@@
  • We propose GMNScore, an outcome reward model that leverages GMN to replace execution-based rewards, achieving both higher efficiency and better performance.
@@
  • We design a novel stepwise reward model StepRTM, which utilizes CTE SQL to deliver stepwise supervision by matching each subquery, resulting in improved accuracy and readability.
@@
  • Extensive experiments show that our reward models consistently improve performance while maintaining high inference efficiency and low GPU memory consumption.
@@ 2 Related Work
Text-to-SQL. Text-to-SQL is a key task in Natural Language Processing (NLP) that involves transforming queries expressed in natural language into executable SQL queries (Tai et al., 2023; Li et al., 2024b; Shi et al., 2025). With the increasing deployment of large language models (LLMs), agentic frameworks (Wang et al., 2025; Pourreza et al., 2025a; Lei et al., 2024) have been introduced to enhance Text-to-SQL tasks. These frameworks enable LLMs to interact with databases through iterative reasoning and external tools. Code Foundation Models such as DeepSeek-Coder (Guo et al., 2024) and Qwen2.5-Coder (Hui et al., 2024) provide the backbone for these agentic systems, enabling structured reasoning and code generation. Several approaches aim to improve LLM performance in Text-to-SQL tasks, including direct fine-tuning (Li et al., 2024a; Yang et al., 2024; Pourreza and Rafiei, 2024), as well as techniques such as prompt design (Pourreza and Rafiei, 2023; Dong et al., 2023; Gao et al., 2024a), self-consistency (Gao et al., 2024a) and schema linking (Guo et al., 2019; Wang et al., 2020; Lei et al., 2020; Lee et al., 2025) to further optimize results.
@@
Reinforcement Learning and Reward Model. RL has become an important paradigm for effectively fine-tuning Code Foundation Models. Policy optimization methods, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024), have been explored. However, the effectiveness of RL training heavily relies on the quality of reward signals, making the design of reward models a critical aspect (Trella et al., 2023). Several contributions to RL-based code generation have advanced reward model strategies. Notable works include CodeRL (Le et al., 2022), which leverages execution feedback; PPOCoder (Shojaee et al., 2023), integrating semantic matching of abstract syntax trees; and AceCoder (Zeng et al., 2025a), applying an LLM-based Bradley-Terry Reward Model.
@@
The execution-based reward model for Text-to-SQL was initially used by Zhong et al. (2017). Recent advancements have introduced continuous reward scores based on keyword matching (Nguyen et al., 2025) and leveraged LLMs to generate reward function candidates and iteratively refine (Berdnyk and Collery, 2025). Alongside these developments, reasoning models such as DeepSeek-R1 (Guo et al., 2025) have advanced RL in reasoning tasks, leading to the introduction of more sophisticated reward model designs. For example, SQL-R1 (Ma et al., 2025) incorporates format and length constraints, while Reasoning-SQL (Pourreza et al., 2025b) employs more complex reward structures, such as schema linking feedback, n-gram similarity scores, and LLM-based judgment. Despite these enhancements, execution-based reward continue to play a central role in the above-mentioned approaches.
@@
Current methods overlook the computational overhead of execution-based and LLM-based reward models and fail to fully exploit the deep semantic structure of SQL queries. Additionally, these approaches focus solely on evaluating the final generated SQL, neglecting the potential of leveraging intermediate supervision signals throughout the SQL generation process. To address these issues, we propose an execution-free outcome reward model and a stepwise reward mechanism. These methods significantly reduce computational overhead while providing more effective reward signals for RL-based Text-to-SQL tasks.
@@ 3.1 Problem Formulation
In the standard Text-to-SQL setting, let $x$ denote a natural language query, $\hat{q}$ and $q^{\star}$ represent the generated SQL and reference SQL query, respectively. In this work, we mainly use Proximal Policy Optimization (PPO) (Schulman et al., 2017), which optimizes the policy model $\pi_{\theta}$ by maximizing:
@@ 3.2 Summary of Existing Reward Models
Recognizing the great importance of reward models in RL, we discuss three types of main reward models. As summarized in Table 1, we compare these models with our proposed reward models in terms of time cost and GPU memory usage during inference. Additionally, the final performance of all reward models is evaluated and ranked, as described in Section 6.1. Detailed information on these comparisons can be found in the Appendix G.
@@
[…]
@@
Table 1: Comparison of Reward Models in RL for Text-to-SQL Tasks. Our proposed GMNScore and StepRTM achieve better performance while significantly reducing time and memory costs.
@@
For the Text-to-SQL tasks, the execution accuracy serves as the most direct reward signal, providing a discrete score based on whether the generated SQL query yields the correct result upon execution. We use a discrete reward model with finer-grained feedback based on syntax error Pourreza et al. (2025b) and runtime diagnostics following Shojaee et al. (2023). Given a generated SQL $\hat{q}$ and reference SQL $q^{\star}$, the formulation is listed as:
@@
However, the EX has notable limitations. When the database contains poor quality data (e.g., limited, missing, or inconsistent entries) or structural issues (e.g., redundancy or anomalies), different queries may produce identical results (Zhong et al., 2020). Test Suite (TS) (Zhong et al., 2020) attempted to address this issue, but as shown in Zhan et al. (2025), false positives and false negatives remain unavoidable. Additionally, repeatedly executing SQL queries introduces significant computational overhead, increasing training time. More details about EX are provided in the Appendix F.
@@
Matching-based Reward.
In (Nguyen et al., 2025), rule-based keyword matching is used for scoring SQL queries, while n-gram similarity is used in Reasoning-SQL (Pourreza et al., 2025b) to capture overlapping token sequences. Matching-based methods are fast and model-free, but may assign negative rewards to semantically equivalent SQL queries that differ in syntax, which should ideally be considered correct. In broader code generation tasks, the PPOCoder (Shojaee et al., 2023) uses semantic matching of abstract syntax trees and data flow graphs. However, it still focuses on surface-level structure and does not fully capture the deep semantic information.
@@ 4 Methodology
We introduce Graph-Reward-SQL, a novel reward model framework designed to enhance SQL generation through two key innovations. First, we propose GMNScore, which replaces EX, reducing time costs while maintaining the accuracy of reward signals without requiring database execution. Second, we introduce StepRTM, a stepwise reward model based on the Relational Operator Tree (ROT) representation of CTE SQL, which provides supplementary intermediate feedback.
Figure 1: Graph-Reward-SQL employs PPO, where rewards drive policy updates. An example illustrates a limitation of EX: the generated SQL ’WHERE age > 34’ and the reference SQL ’WHERE age >= 34’ produce identical results despite their semantic difference. In contrast, our proposed GMNScore leverages graph representation to capture deep semantic similarity. Additionally, a stepwise reward model, StepRTM, that is tailored for CTE SQL addresses the lack of intermediate rewards. A mock stepwise score is provided for illustration; see Figure 2 for more details.
@@ 4.1 Relational Operator Tree (ROT)
To bridge this gap, we leverage the Relational Operator Tree (ROT) to represent SQL queries as trees of relational algebra operators. Each node in the tree corresponds to a specific logical operation (e.g., Join, Project, Filter), while the tree structure itself reflects the dependencies and execution order of the query. In practice, we use Apache Calcite (Begoli et al., 2018) to generate ROTs, which compiles SQL into a canonical intermediate representation called RelNode. This format includes various optimizations, such as operator reordering and clause simplification, resulting in normalized logical plans that are more resilient to surface-level differences. Similar to CFGs and DFGs, the RelNode format can also integrate control dependencies and data flow as edges (Zhan et al., 2025). This enables the creation of more comprehensive graph representations that include richer SQL semantics, facilitating query understanding and evaluation.
@@ 4.2 FuncEvalGMN
After obtaining the SQL graph representations $G_{1}$ and $G_{2}$, we employ a Graph Matching Network (GMN) (Li et al., 2019) trained in SQL pairs (Zhan et al., 2025) to assess functional equivalence. It is trained using contrastive learning for pretraining and supervised learning to capture deep semantic similarity of SQL queries. The similarity between two queries is computed as the negative Euclidean distance between their final graph-level embeddings: $s(h_{G_{1}},h_{G_{2}})=\big\lVert h_{G_{1}}-h_{G_{2}}\big\rVert_{2},$ where $h_{G_{1}}$ and $h_{G_{2}}$ are computed by the GMN, considering the joint representations of $G_{1}$ and $G_{2}$. This approach, first introduced in FuncEvalGMN (Zhan et al., 2025), is described in further detail in Appendix O, and the details of our further optimization are provided in Appendix E.